OpenStudy (anonymous):

what is the probability that 19 tosses of a fair coin will show 8 heads

OpenStudy (ibbutibbu.):

i would solve this using proportions: \[\frac{ 8 }{ 19 } = \frac{x}{100}\]

OpenStudy (midhun.madhu1987):

The value of probability always lies between 0 and 1 @Ibbutibbu.

OpenStudy (ibbutibbu.):

i know, onece you get the \frac{x}{100} nad find out what x is you can turn it in to a percent

OpenStudy (midhun.madhu1987):

is it 8/38?

OpenStudy (ibbutibbu.):

in this case x is approximately 42. then .42, which is approximatesly 42% @midhun.madhu1987

OpenStudy (ibbutibbu.):

but, you may be right

OpenStudy (anonymous):

0.0288 0.0721 0.1442 0.2884 these are the options

OpenStudy (ibbutibbu.):

@midhun.madhu1987 , what do you think?

OpenStudy (midhun.madhu1987):

still thinking.. :)

OpenStudy (ibbutibbu.):

Hard one, huh?

OpenStudy (anonymous):

19C8*(1/2)^8*(1/2)^11... binomial theorem :)

OpenStudy (anonymous):

just figured it out... thanks guys